• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name MinaKat

    Log entry time 00:04:55 on December20,2007

    Entry number 222311

    keyword=Shift Summary

    Shift Summary:
    ===============

    4:15 Finished BCM calibrations (started in previous shift)

    We proceed with H2/He target boiling test (part #2)

    Asked for beam raster 2x2, set at BPM A and B at -2mm
    Momentum set (both arms): 2.791 GeV/c, angle (both arms): 15 deg

    Target: He Loop 1, raster 2x2
    -------------------------------
    Run (Right/Left) Current(uA) Events (Right/Left)
    ------------------------------------------------------
    22173/4317 5
    22174/4318 10 0.7M/0.3M
    22175/4319 20 1.2M/0.7M
    22176/4320 30 1.4M/0.6M
    22177/4321 40 0.9M/0.5M
    22178/4322 50 1.2M/0.6M
    22180/4324 60 1.1M/1.1M
    22181/4325 70 1.1M/1.1M



    Left Spectrometer problem: when quads were cycled, regulator was not
    set back for Runs 22173 - 22178. Boiling tests are ok for the Right
    spectrometer. Set regulator on.
    1
    22179/4322 test.

    Target: C12 Left, ratster 2x2
    -----------------------------
    Run (Right/Left) Current (uA) Events (Right/left)
    ---------------------------------------------------------
    22182/4326 5 0.1M/0.1M
    22183/4327 10 0.3M/0.3M
    22184/4328 20 0.5M/0.5M
    22185/4329 30 0.5M/0.4M
    22186/4330 40 0.7M/0.7M
    22187/4331 50 0.8M/0.8M
    22188/4332 60 0.5M/0.5M
    22189/4333 70 0.6M/0.6M

    Target: He Loop 1, raster size 3x2
    ----------------------------------

    Run 22190 test raster size

    Run (Right/Left) Current (uA) Events (Right/Left)
    --------------------------------------------------------
    22191/4335 5 0.4M/0.4M
    22192/4336 10 0.8M/0.8M
    22193/4337 20 1.0M/1.0M
    22194/4338 30 1.0M/1.0M
    22195/4339 40 1.1M/1.1M
    22196/4340 50 1.0M/1.0M
    22197/4341 60 1.0M/1.0M
    22198/4342 70 1.0M/1.0M

    Momentum set (both arms):2.573 GeV/c

    Target: H2 Loop 3, raster size 2x2
    ----------------------------------
    Run 4343 test raster size 5uA (good run)

    20:44 While right spectrometer dipole is setting, Alex will do
    NaI HV balancing. So put optics target in.
    Runs 4344,4345,4346 HV tests

    21:10 Dipole finally settles, put H2 Loop 3 target back in

    Target: H2 Loop 3 raster size 2x2
    ---------------------------------
    Run (Right/Left) Current (uA) Events (Right/Left)
    ------------------------------------------------------------------
    22199/4347 5 Right/Left spect. IOC die



    21:16 Right IOC dies. Rebooted (during 5uA runs 22199/4347), Dipole NMR will not lock. Next we loose Left IOC.....Cannot reboot from hareboot10

    Paged tech on call,
    Tech is on his way, asked for hall entry
    1 hour later right dipole NMR finally locks.

    22:40 Tod in hall
    J.P.'s instructions:
    1) scratch remaining boiling test points.
    2) proceed with item 3 (Optics) on Runplan.
    3) run with the right spectrometer only, if IOC
    problems persist.
    23:45 Tech still in hall, Alex goes in too.

    00:01 Alex calls from hall. Network switch was down.
    We just got back the left spectrometer display in Hall A tools.!!!!!